OpenStudy (anonymous):

the slope of a line is -1/3. The point (-12, -6) is on the line. Determine the equation of this line

OpenStudy (lalaly):

y-(-6)= -1/3 (x--12) y+6=-1/3 x -4 y= -1/3 x -10